ayomide40:

any one taking university physics A particle is moving around in a circle and its position is given in polar coordinates as x = Rcosθ, and y = Rsinθ, where R is the radius of the circle, and θ is in radians. From these equations derive the equation for centripetal acceleration.

Rx1nyDxyz:

does physical science count or no?

ayomide40:

not really

Rx1nyDxyz:

Then im not sure i can help, but ill try. What do u need help on?

Ashely:

It depends on what you need help on

Rival:

Do you have any school-related questions??

ayomide40:

A particle is moving around in a circle and its position is given in polar coordinates as x = Rcosθ, and y = Rsinθ, where R is the radius of the circle, and θ is in radians. From these equations derive the equation for centripetal acceleration.

Rx1nyDxyz:

Ok, i think i can help.

ayomide40:

kk

Rival:

@ayomide40 wrote:
A particle is moving around in a circle and its position is given in polar coordinates as x = Rcosθ, and y = Rsinθ, where R is the radius of the circle, and θ is in radians. From these equations derive the equation for centripetal acceleration.
Could you put that as your main question (edi)?

ayomide40:

wait is this the right answer d2x/dt2 = -Rw2

Rx1nyDxyz:

We only need to look at the equation for the x-position, as we know that centripetal acceleration points towards the center of the circle.

Ashely:

do you have a ss of this all being seperated apart

Rx1nyDxyz:

So, when θ = 0, the second derivative of x with respect to time has to be the centripetal acceleration.

Rx1nyDxyz:

So, what do you think the first derivative of x is?

ayomide40:

dx/dt = -Rsinθ(dθ/dt)

Rx1nyDxyz:

Correct, and what you think the second is?

ayomide40:

d2x/dt2 = -Rcosθ(dθ/dt)2−Rsinθ(d2θ/dt2)

Rx1nyDxyz:

That is correct, because we usethe chain rule of Calculus and by assumption θ is a function of time.

Rx1nyDxyz:

So, θ can be differentiated with respect to time. Now we have to evaluate the second derivative at θ = 0.

Rx1nyDxyz:

Do you know wat comes next?

ayomide40:

is it this am not sure d2x/dt2 = -R(dθ/dt)2

Rx1nyDxyz:

You're close.

Rx1nyDxyz:

dθ/dt is called the angular velocity, which is the rate of change of the angle θ. So, we'll let w ≡ dθ/dt.

Rx1nyDxyz:

So, your final answer is d^2x/dt2 = -Rw^2

ayomide40:

kk ty

Rx1nyDxyz:

yw anyday

atoxty59:

gj